### v1

# Genie Sim 3.0 : A High-Fidelity Comprehensive Simulation Platform for Humanoid Robot

Abstract
The development of robust and generalizable robot learning models is critically contingent upon the availability of large-scale, diverse training data and reliable evaluation benchmarks. Collecting data in the physical world poses prohibitive costs and scalability challenges, and prevailing simulation benchmarks frequently suffer from fragmentation, narrow scope, or insufficient fidelity to enable effective sim-to-real transfer. To address these challenges, we introduce Genie Sim 3.0, a unified simulation platform for robotic manipulation. We present Genie Sim Generator, a large language model (LLM)-powered tool that constructs high-fidelity scenes from natural language instructions. Its principal strength resides in rapid and multi-dimensional generalization, facilitating the synthesis of diverse environments to support scalable data collection and robust policy evaluation. We introduce the first benchmark that pioneers the application of LLM for automated evaluation. It leverages LLM to mass-generate evaluation scenarios and employs Vision-Language Model (VLM) to establish an automated assessment pipeline. We also release an open-source dataset comprising more than 10,000 hours of synthetic data across over 200 tasks. Through systematic experimentation, we validate the robust zero-shot sim-to-real transfer capability of our open-source dataset, demonstrating that synthetic data can server as an effective substitute for real-world data under controlled conditions for scalable policy training. For code and dataset details, please refer to: https://github.com/AgibotTech/genie_sim.
I INTRODUCTION
The advancement of robotic manipulation is increasingly supported by vision-language-action (VLA) models, which equip robots with the ability to interpret natural language instructions and execute corresponding physical actions [undef] [undefa]. This capability allows robots to perform a wider range of tasks in unstructured environments, moving closer to human-like adaptive operation [undefb]. However, the development of robust and generalizable VLA models relies heavily on access to large-scale, high-quality datasets of robot interactions, which are resource-intensive and time-consuming to collect in the real physical world [undefc][undefd]. Furthermore, systematically evaluating and iterating these models requires a consistent, reproducible and scalable testing environment, but the current evaluation approaches on real devices often fail to meet this requirement [undefe].
A high-fidelity simulation platform is a practical and efficient approach to provide large-scale, cost-effective and low sim-to-real gap synthetic dataset and benchmark for VLA training and evaluation [undeff][undefg]. A well-designed simulation platform can provide the necessary scale, diversity, and repeatability for both training and evaluation, significantly accelerating the development cycle of VLA-driven robotic systems [undefh][undefi]. However, scaling this process encounters three fundamental bottlenecks. First, creating high-fidelity simulation environments that accurately reflect the complexity of the real world traditionally requires significant manual efforts from specialists in 3D modeling and physics simulation. This process is time-consuming, labor-intensive and limits the scale and variety of training data, becoming a critical barrier to model generalization [undefj][undefk][undefl]. Second, while automated or procedural generation can create varied scenes, it often lacks fine-grained control. A core tension exists between generating sufficient diversity for robust learning and maintaining the ability to precisely reproduce or systematically vary specific scenarios. This makes it difficult to debug failures, conduct controlled ablation studies, or measure systematic generalization—hindering rigorous model development [undefm]. Third, current model evaluation relying on a fixed set of handcrafted metrics (e.g., success rate) fails to capture nuanced task completion quality[undefn]. Furthermore, human-in-the-loop evaluation is inefficient, subjective, and non-scalable, resulting in an incomplete assessment of policy performance and slows down the iterative training-evaluation cycle.
To resolve above-mentioned issues, we introduce Genie Sim 3.0, a comprehensive open-source simulation platform for environment reconstruction, scene generalization, data collection and automated evaluation. Genie Sim 3.0 integrates the following four novel features: (1) large language model (LLM)-driven scene generation and generalization; (2) LLM-based task and evaluation generation; (3) high fidelity simulation via 3D reconstruction and visual generative synthesis; (4) dual-mode data collection pipeline, including teleoperation and automation.
Developing on these innovative new features, Genie Sim 3.0 provides five new simulation modules to significantly accelerate the model development of embodied intelligence: (1) Genie Sim asset retrieval system, supporting sematic query-based search for 5,140 simulation-ready objects across 353 categories; (2) a scene generation pipeline that operates solely on natural language input and is equipped with comprehensive domain randomization capabilities; (3) more than 10,000 hours of synthetic data with multi-dimensional domain randomization across 200 tasks; (4) a comprehensive benchmark integrating close-loop evaluation with over 100,000 evaluation scenarios, designed to assess a holistic capability profile across multiple dimensions, covering semantic understanding, spatial reasoning, and operational execution, therefore clearly defining model’s performance boundaries and optimization directions.
Contributions of Genie Sim 3.0 can be summarized into five key aspects:
  1. 1.
    We develop a scene generation pipeline called Genie Sim Generator that features a natural language interface that interprets high-level semantic instructions to generate highly realistic simulation scenes in real-time. This interface supports multi-round conservation for iterative refinement, significantly improving efficiency of scene generation.
  2. 2.
    We support rapid, multi-dimensional scene generalization from a single generated scene mentioned above. It can produce vast and varied scenarios by parameterizing factors such as lighting, background, layout, poses, trajectories, sensor noise, and robot morphology within several minutes via natural language instructions.
  3. 3.
    We establish a scalable evaluation benchmark built upon more than 100,000 simulation scenarios. It leverages LLM to automatically generate task instructions and evaluation protocols, which are then executed and scored by VLM. This constructs a comprehensive and multi-dimensional capability profile—spanning semantic understanding, spatial reasoning, and operational execution—that clearly delineates a model’s performance boundaries and optimization pathways.
  4. 4.
    Through systematic experiments, we validate the effective sim-to-real transfer capability of the proposed benchmark framework and synthetic data. Empirical results demonstrate that, under certain conditions, the generated synthetic data can serve as a viable substitute for real-world robotic data, thereby providing scalable support for model training. Furthermore, the benchmark enables the effective prediction and extrapolation of a model’s performance boundaries in real-world scenarios based on its performance in simulation, offering stable and reproducible guidance for model iteration.
  5. 5.
    We open-source Genie Sim 3.0 simulation platform which integrates scene generation, evaluation synthesis, data acquisition, and closed-loop model assessment. To foster community progress, we open-source 5140 assets, 10,000+ hours dataset, 100,000+ evaluation scenarios, and the complete codebase. By providing massive-scale synthetic dataset and evaluation scenes for insightful capability diagnostics and panoramic evaluation to accelerate model iteration, Genie Sim 3.0 reduces reliance on physical hardware, enhances research and development efficiency, and propels the development of practical embodied AI applications.
II RELATED WORK
II-A Robotics Dataset
Numerous open-source datasets are available in robotics, encompassing data collected from both physical robotic platforms and simluation environments. DROID [undefo] dataset provides over 70,000 real-world robotic manipulation demonstrations; RoboNet [undefp] integrated 160,000 trajectory data from 7 different robot systems; RH20T [undefq] includes 100,000 demonstrations spanning 147 different tasks, reflecting a high degree of task diversity; Open X-Embodiment [undefr] integrates vast amounts of heterogeneous data from laboratories worldwide, comprising over 1 million trajectories, 160,266 tasks, and 22 distinct robot embodiments; Agibot World [undefs] comprises millions of real-robot data trajectories, covering five core scenarios and encompassing over 80 skills. Despite the extensive scale, real-world datasets impose two fundamental limitations: prohibitively high acquisition costs and narrow generalization capability.
Synthetic datasets provide cost-effective, large-scale, diverse data with controlled domain randomization to promote robust policy generalization. RoboCasa [undeft] offers more than 100,000 trajectories, yet its scenes are predominantly confined to kitchen settings; DexGraspNet 2 [undefu] contains over 400 million demonstrations but focuses solely on grasping actions without dynamic scene variations or more complex manipulation skills; RoboTwin 2.0 [undefv] encompasses over 100,000 bimanual manipulation trajectories across 50 tasks, but the physical and visual fidelity of its simulation environment remains notably constrained.
II-B Robotics Benchmark
Benchmarks are essential for assessing model performance, yet creating high-quality benchmarks continues to be a significant challenge with the field. Meta-World [undefw] provides a benchmark designed for multi-task and meta-reinforcement learning. HumanoidGen [undefx] establishes a dedicated benchmark comprising 20 tabletop manipulation tasks. HumanoidBench [undefy] contains 27 whole-body control tasks designed to evaluate reinforcement learning and hierarchical control methods. Bigym [undefz] introduces a benchmark comprising 40 manipulation tasks set in household kitchen scenarios. BEHAVIOR-1K [undefaa] provides a comprehensive benchmark that includes 50 diverse scenes, over 9,000 object models spanning 1,900+ categories. ManipulaTHOR [undefab] is a benchmark consisting of 30 kitchen scenes, over 150 object categories, with defined training, validation and test splits and evaluation metrics. Although seminal benchmarks such as HomeRobot [undefac], DaXBench [undefad], SoftGym [undefae], PlasticineLab [undefaf], INT-ACT [undefag], ManiFeel [undefah], AGNOSTOS [undefai], RoboTwin 2 [undefv], VLABench [undefaj] contribute valuable evaluations, the field remains constrained by several persistent challenges, including narrow evaluation scope, notable sim-to-real gaps in both physical dynamics and visual rendering, and insufficient transfer efficacy to real-world deployments.
III Method
III-A Scene Generation
Figure 2: The Automated Workflow of Genie Sim Generator. This module captures user intent via multi-round conversation, translates it into executable Python code, and compiles the final scene graph with assets for Isaac Sim.
With the proliferation of LLM and VLM, scalable, generalizable and conversational data creation has become an emerging paradigm.
We implement Genie Sim Generator, an LLM-driven toolchain that jointly creates simulation scenes and produces Scene Graph for downstream task evaluations. It consists of two tightly-coupled modules. The first is Assets Index, a managed repository of Isaac Sim-ready assets augmented with LLM, VLM and retrieval-augmented generation (RAG) agents, which offers structured Application Programming Interface (API) to facilitate scene composition. The second is Scene Generator, inspired by the scene language [undefaj], this module captures user intent via multi-round conversation, translates it into executable Python code, and compiles the final scene graph with assets for Isaac Sim.
The entire scene generation process has four stages: intention interpret, assets retrieve, domain specific language (DSL) code generation and results assemble, all within the same chat window sharing the same context.
Intention Interpreter translates user’s open-ended natural language prompt into a structured, machine-interpretable task request. A chain-of-thought (CoT)–enabled LLM first parses and decomposes the natural language prompt into a structured decomposition comprising spatial scene descriptors, object attribute constraints and task-level intents (e.g., “stack-up”, “tidy”, “random”). Ambiguous or underspecified phrases (e.g., “some blocks”, “randomly arranged”) are resolved through a reasoning process grounded in a pre-trained world-knowledge memory. The resulting specification is formulated as a JavaScript Object Notation (JSON) schema containing required semantic object classes (with optional geometric constraints such as size, color and shape) and pairwise spatial relations (e.g., “on”, “adjacent”, “aligned”). The resulting schema is forwarded to the Assets Index to initiate retrieval process. If constraints contradict predefined system rules, the interpreter engages an explicit feedback loop with the user for clarification.
Assets Index is a RAG-powered asset retrieval module. Semantic descriptions are first extracted for all 5,140 objects based on their appearance, geometry, and usage. Subsequently, these descriptions are encoded into 2048-dimensional vectors via the QWEN: text-embedding-v4 model and stored in a ChromaDB vector database. At runtime, the planner extracts keywords (e.g., “yellow cube”) from scene description and encodes them into the same embedding space using the identical model. The resulting query vector is compared against the stored asset embeddings via cosine similarity, and the top-k most similar candidates are retrieved along with their metadata—including Universal Scene Description (USD) paths, collision hulls, mass properties, and texture variants. These assets are seamlessly integrated into the LLM context, ensuring that subsequent code generation references only pre-validated, available assets. This retrieval step is fully transparent to the user and typically completes within 200 milliseconds.
DSL Code Generater draws upon the syntactic structure defined in the scene language [undefaj], but extends it with a back-end adaptation for seamless interoperability with Genie Sim Assets Library. Integrating the contexts from Intention Interpreter, Assets Index and DSL definition, our system synthesizes a precise scene specification via pre-trained LLM. This representation exhibits fine-grained controllability, double float precision and enhanced generalization capability. We maintain the chat context within our engine to support iterative task editing. The output can also be manually adjusted to address limitations of LLM. By embedding our asset library into the LLM context, we achieve joint generalization over categories, poses, lighting, and textures without any fine-tuning. Rich semantic annotations and geometric properties within assets library enables scene-level generalization. Complicated layouts which extend beyond basic tabletop scenarios are also supported such as shelves and storage racks.
Results Assembler finally collects the output from chat context and instantiates the DSL program generated by LLM. To achieve randomization, the DSL program employs random functions that introduce variability in object poses, layout patterns, and choice of objects. During instantiation, a hierarchical Scene Graph with two major parts is created: nodes (objectes encoded with asset id, semantic, size, pose and task tag) and edges (spatial relations such as on, in, adjacent, aligned and stacked). Finally assembler utilizes the OpenUSD Schema and Isaac Sim APIs to synthesize simulation-ready USD files. Our Results Assembler generates thousands of diverse scenes within minutes, leveraging the power of LLM and advanced algorithms to produce massive scene layouts with high efficiency.
Figure 3: VLM-Driven Evaluation.
III-B Evaluation Generation
Current open-source simulation benchmarks typically rely on predefined instructions, manually annotated success criteria and repeated trial executions to evaluate VLA models. This paradigm yields a largely unidimensional instruction space, limits the scalability of evaluations, and entails high costs for new task creation. LLM exhibits strong natural language understanding and therefore can be introduced for automated instruction generation, task decomposition and prompt expansion, enabling more diverse, scalable, and cost‑effective evaluation pipelines.
Given a certain simulation scenario, Genie Sim Benchmark leverages the capabilities of the LLM by combining it with Action Domain Evaluation Rule (ADER) system to automatically generate numerous reasonable instructions and executable evaluation configuration files, which are then integrated with the simulator for automated evaluation.
VLM demonstrates strong visual and semantic understanding. Given formalized task specifications alongside the temporal sequence of visual observations recorded during task execution, VLM can efficiently determine whether task requirements have been satisfied and generate evidence‑based justifications (see Fig. 3) . This evaluation paradigm is applicable to the majority of simulation benchmark tasks, markedly reducing human annotation effort and allowing the creation of large‑scale evaluation instances at low cost.
Figure 4: Automated Data Collection. A complete task parsing and execution pipeline, which improves task success rate through waypoint filtering and a robust retry mechanism.
III-C Environment Reconstruction
To provide large-scale, high-fidelity interactive simulation environment, we utilize 3DGS [undefak] neural rendering technology for photorealistic image rendering and surface reconstruction, generating high-precision meshes for interactive physical simulation. We collect data using SkylandX Innovation’s MetaCam handheld 3D laser scanner. The collected data contains fisheye images, the corresponding poses for each frame, and the dense point cloud data of the entire scene.
However, 3D Gaussian Splatting (3DGS) imposes relatively stringent requirements on camera pose accuracy; deviations of even a few pixels can result in blurred rendering, artifacts, and geometric bulges. In complex indoor environments, LiDAR‑based SLAM often cannot achieve pixel‑level positioning precision, leading to insufficient pose accuracy for high‑quality 3DGS reconstruction. In the camera pose optimization module, we first utilize SuperPoint [undefal] and LightGLue [undefam] to replace the DSP-SIFT [undefan] and feature extraction module in COLMAP-PCD [undefao], which reduces noise in 2D feature points and enhances the ability to extract features in weak textures. Then, we perform triangulation directly using the prior poses of the camera obtained by LiDAR SLAM, and search the 2D feature points corresponding to the lidar 3D points for association and the sparse 3D points to implement BA optimization together. After obtaining the camera pose and 3D point data output using COLMAP-PCD, we train 3DGS using the open-source gsplat [undefap]framework. During the process of large-scale scene reconstruction, it is always inevitable to collect sufficient coverage of the views. To compensate for the insufficient number of views during the collection process and improve the rendering quality, we utilize generative models to extrapolate sufficient views. We use the pretrained model from Difix3D+ [undefaq] to render extrapolated views and obtain high-quality images. Finally, we conducte 3DGS training based on surface reconstruction using a large number of new perspective images produced by the diffusion model and the corresponding pose data, as well as the lidar point cloud data obtained during data collection. We use PGSR [undefar] for surface reconstruction and obtained a high-precision mesh.
III-D Data Generation
The data collection framework integrates two complementary paradigms: teleoperation and automation. Teleoperation is utilized for complex long-horizon tasks to produce high quality human-like demonstrations. In contrast, automated collection excels in efficiency and cost-effectiveness, facilitating the rapid generation of extensive synthetic data for robot imitation learning.
Teleoperation: Our teleoperation framework utilizes a PICO VR Head-Mounted Display (PICO) device to bridge human input with a simulated robotic environment. Functioning as the primary input apparatus, the PICO sends action signals representing the target end-effector pose to a central host. A benchmark module processes these signals, and a motion controller executes the planned trajectory, driving the virtual motion of the robot in simulation environment.
The simulation system incorporates physical effects in real world, including collisions and friction, which enables the robot to manipulate and interactive with simulated objects. The entire interaction sequence, encompassing robot joint states, visual observations, and object poses, is fully logged to constitute the synthetic dataset. Consequently, this approach generates anthropomorphic motion data and leverages human expertise to tackle intricate tasks that are difficult to automate.
Automated Collection: Leveraging cuRobo [undefas]—a GPU-accelerated motion planner—as the core trajectory planning module, the fully automated data collection pipeline consists of two parts: task generation and data collection (Figure 4). Task generation is facilitated by an LLM-based asset retrieval system, through which required simulation assets can be efficiently queried and assembled using predefined atomic skills.
During task parsing, candidate key waypoints—derived from pre-annotated asset configurations, such as grasping poses labeled by GraspNet [undefat]—are evaluated based on kinematic reachability, collision avoidance, and anthropomorphic feasibility. To improve robustness, multiple candidate waypoints are generated per action, forming several alternative action sequences. Each sequence is executed in simulation and subsequently assessed by a dedicated trajectory evaluation module. In case of execution failure, a state rollback is performed before the next candidate sequence is attempted.
Unlike most previous approaches that simplify the planning environment—for example, objects not directly related to the task are excluded, which can lead to collisions in cluttered scenes—our system retains environmental completeness. In order to balance completeness and efficiency, mesh simplification is applied to object geometries during scene initialization, maintaining trajectory reliability while significantly improving computational efficiency during planning.
III-E Closed-loop Evaluation
For model closed-loop evaluation, the simulation and the model inference environment are decoupled and communicate over the Hypertext Transfer Protocol (HTTP) protocol. The simulator transmits the robot’s observation images and proprioceptive states to the inference service and the model returns control commands, which are then executed in the simulation environment. During simulation execution, task completion is periodically evaluated; if the task is completed, the simulation terminates, otherwise a predefined timeout procedure is invoked to terminate the run. Genie Sim 3.0 benchmark supports the following evaluation features:
  • Integration with common VLA models ($\pi_{0.5}$ [undefau], GO-1 [undefs], UniVLA [undefav], RDT [undefaw], X-VLA [undefax], etc.)
  • Multiple robot types (Genie G1, Genie G2)
  • Multiple end effectors (omnipicker, omnihands, INSPIRE skillhands, zhixing gripper, etc.)
  • Local/distributed inference framework
  • Automated and multi-dimensional evaluation
IV Dataset
IV-A Task Distribution
To accommodate developers at different stages who may prioritize varying levels of data complexity and focus, our dataset is designed to provide comprehensive coverage of the core competencies required by contemporary VLA models for embodied intelligence. The complexity of the tasks is hierarchically organized, facilitating a structured progression from simple to complex scenarios. A key design principle is composability: long-horizon tasks can be decomposed into sequences of fundamental sub-tasks contained within the dataset.
To this end, we structure the task taxonomy along three primary axes (Figure 5): manipulation skill, cognitive comprehension, and task complexity.
Figure 5: Task Distribution Matrix. The dataset is constructed along three dimensions: Manipulation skill, Cognitive comprehension, and Task complexity.
  • Atomic Skills encompass the model’s fundamental motor actions, such as pick, place, pull, push, open, and close.
  • Cognitive Comprehension pertain to the model’s ability to interpret tasks, scenes, and instructions. This includes competencies like spatial reasoning (e.g., size relationships), attribute understanding (e.g., color), logical inference, and commonsense reasoning.
  • Task Complexity is defined along metrics such as planning horizon and the need for coordinated control. For instance, the sequence from “single-arm removal of a trash item” to “bimanual coordinated removal of a trash item” and finally to “cleaning all trash from a desktop” illustrates a graduated increase in complexity.
IV-B Data Distribution
Building upon the aforementioned principles of task distribution, we introduce Genie Sim 3.0 synthetic dataset. This dataset encompasses 200 representative tasks for embodied intelligence, providing over 10,000 hours of simulated interaction that reflects its substantial scale. The data is generated using two robotic platforms, Agibot’s G1 and G2, and incorporates systematic variations across multiple dimensions—including task layout, initial robot pose, environmental lighting, scene configuration, camera noise, and semantic instruction phrasing—ensuring broad coverage and strong generalization capability. The resulting dataset thus supports the training and evaluation of embodied intelligence models with high-quality, diverse, and robust simulated experiences.
V Experiment
TABLE I: Experiment Results: four tasks, each with four models tested in both real and sim environments
Select Color Recognize Size Grasp Targets Organize Objects
sim env real env sim env real env sim env real env sim env real env
200 eps real 0.45 0.53 0.50 0.56 0.34 0.39 0.25 0.30
500 eps real 0.75 0.73 0.75 0.75 0.54 0.58 0.45 0.40
500 eps sim 0.53 0.60 0.50 0.63 0.29 0.33 0.39 0.35
1500 eps sim 0.86 0.85 0.93 0.94 0.80 0.71 0.52 0.60
The validity and reliability of simulation benchmarks and datasets constitute a foundational consideration in embodied intelligence research.
Figure 6: Sim to Real Experiments. Comparison of initial task setup between real and sim testing environments
Two central questions are examined: first, whether a simulation benchmark can evaluate model performance as reliably as physical environment; and second, to what extent data synthetic data can effectively substitute for physically-collected data in model training. To address these questions systematically, a series of experiments are designed to assess the validity and efficacy of Genie Sim 3.0 benchmark and dataset.
V-A Experimental Setup
In this work, we conducted 32 groups of experiments to compare the task success rates of models fine-tuned with varying amounts of real-world and synthetic data, tested in both physics and simulation environments. The $\pi_{0.5}$ model [undefau]— a widely recognized baseline with demonstrated performance in related research— is selected as the base model and Agibot G1 robot is deployed for task execution.
Since the primary objective of these experiments is to evaluate the validity and reliability of the simulation benchmark and dataset, tasks with either excessively high or low success rates are avoided to better reveal performance differences across dataset and environments. In accordance with this principle and to ensure assessment covering both atomic skills and cognitive capabilities, four representative tasks are selected for evaluation:
  1. 1.
    Select Color,
  2. 2.
    Recognize Size,
  3. 3.
    Grasp Targets,
  4. 4.
    Organize Items.
For each task, models are post-trained with four different setups:
  1. 1.
    200 episodes of real-env data,
  2. 2.
    500 episodes of real-env data,
  3. 3.
    500 episodes of sim-env data,
  4. 4.
    1500 episodes of sim-env data.
Each experiment is evaluated over 50 trials in real world environment and over 250 trails in simulation.
Figure 7: Model Performance on Real-World Environment. The success rates of the 4 tasks, each with four training setups are shown in the figure with different colors.
V-B Results and Analysis
Figure 8: Performance Comparison between Real and Sim Environments. The performance in simulation share the same trend with it in real-world environment.
Effectiveness of the Synthetic Data. In general, the success rates of the tasks significantly improve with an increasing volume of training data—for both real and synthetic data—which aligns well with the scaling laws in learning performance, see Fig. 7. Increased task complexity leads to reduced success rates for models trained on datasets of identical size. Most importantly, the model trained exclusively on 1500 eps of synthetic data attains the highest zero‑shot success rates across all four tasks in the real‑world environment, as shown in Tab. I. When comparing data of the same quantity (500 eps), models trained on real‑robot data consistently outperform those trained on synthetic data. This advantage can be attributed to the higher physical fidelity of real data, which inherently captures complex real‑world phenomena such as friction, contact dynamics, and collisions—aspects that are often imperfectly modeled and lead to sim‑to‑real gaps in synthetic datasets.
Figure 9: Correlation Analysis of Model Performance on Sim and Real Environments. All 16 models are evaluated both in sim and real environments, which are tagged on the right side of the figure, each with a different color and marker.
Figure 10: Success Rate on both Sim and Real Env in Heatmap. The comparison of the left and right charts intuitively reflects the consistency between the real and the sim environment.
Nevertheless, real‑world robotics data collection incurs substantial costs and time constraints, limiting the scale and diversity of available datasets. In contrast, synthetic data can be generated efficiently at scale, allowing for extensive randomization in parameters such as object textures, lighting conditions, physics properties, and task variations. By scaling the synthetic dataset to 1500 episodes and systematically varying simulation conditions, the resulting training distribution effectively mitigates the domain gap, thereby enhancing the model’s generalization capabilities to real‑world settings.
Thus, the large‑scale, systematically randomized synthetic data pipeline implemented in GenieSim 3.0 offers a pragmatic and scalable alternative to real‑robot data collection, enabling the training of robust models capable of high zero‑shot performance in physical environments.
Consistency between Sim and Real Environments. Fig. 8 provides a comparative analysis of model performance in simulated versus physical environments. While minor discrepancies are observed across tasks—primarily attributable to systematic differences from approximations of contact dynamics and friction in simulation, as well as stochastic factors inherent to real-world deployment such as unstructured environmental perturbations—the overall performance trends remain consistent. This close correspondence is further illustrated in the success-rate heatmaps presented in the Fig. 10. A quantitative correlation analysis, summarized in the scatter plot of the Fig. 9, reveals a statistically significant strong positive relationship between simulated and real-world performance ($R^{2}=0.924$), with a best-fit slope $\approx 1.045$. These results indicate that evaluations conducted in Genie Sim 3.0 yield reliable conclusions that are consistent with trends observed in real-world testing.
Given this high level of alignment, alongside the advantages of simulation—including fully automated evaluation pipelines, high reproducibility, scalable scenario generation, support for concurrent batch executions, and significantly lower operational costs—Genie Sim 3.0 serves as a robust and efficient platform for model assessment. Its use can substantially accelerate research iteration and facilitate experimental scaling beyond the practical constraints of physical testing.
VI CONCLUSIONS
We present Genie Sim 3.0, a unified platform that addresses key bottlenecks in robot learning through these integrated contributions: Genie Sim Generator that uses LLM to create generalized, high-fidelity scenes from natural language instructions; a pioneering LLM-VLM benchmark for automated, scalable task generation and evaluation; and a large-scale open-source dataset validated for effective zero-shot sim-to-real transfer. Together, these components constitute a cohesive foundation for advancing scalable data generation, robust policy assessment, and reproducible research toward more generalizable robotic manipulation.
APPENDIX
ACKNOWLEDGMENT
We would like to express our sincere gratitude to Sukai Wang and Xiaowei Cai for their assistance with model inference and training; to Haowen Yin, Yanping Zhou, Zengcheng Zhou, and Cheng Ruan for their support in experimental testing; to Chengyue Zhao, Xinke Yu, and Jialu Li for their work on publicity and design; and to Cheng Jing, Haoyu Cao, Chi Zhang, Guangte Xiang, Jun Dai for their help with data processing.

### v2

@@
     We open-source Genie Sim 3.0 simulation platform which integrates scene generation, evaluation synthesis, data acquisition, and closed-loop model assessment.
-To foster community progress, we open-source 5140 assets, 10,000+ hours dataset, 100,000+ evaluation scenarios, and the complete codebase.
+To foster community progress, we open-source 5,140 assets, 10,000+ hours dataset, 100,000+ evaluation scenarios, and the complete codebase.
 By providing massive-scale synthetic dataset and evaluation scenes for insightful capability diagnostics and panoramic evaluation to accelerate model iteration, Genie Sim 3.0 reduces reliance on physical hardware, enhances research and development efficiency, and propels the development of practical embodied AI applications.
@@
 Genie Sim 3.0 benchmark supports the following evaluation features:
-  • Integration with common VLA models ($\pi_{0.5}$ [undefau], GO-1 [undefs], UniVLA [undefav], RDT [undefaw], X-VLA [undefax], etc.)
+  • Integration with common VLA models ($\pi_{0.5}$ [undefau], GO-1 [undefs], GR00T, UniVLA [undefav], RDT [undefaw], X-VLA [undefax], etc.)
   • Multiple robot types (Genie G1, Genie G2)
@@
 V Experiment
-TABLE I: Experiment Results: four tasks, each with four models tested in both real and sim environments
+To systematically assess the validity and utility of Genie Sim as a simulation benchmark and data generation platform, we introduce four closed-loop evaluation suites.
+GenieSim-Sim2Real is designed to validate the data quality of Genie Sim and the fidelity of its simulated environments: it examines whether large-scale synthetic data can match or exceed real-world data in model training, and whether benchmark evaluations in simulation yield conclusions consistent with physical deployment.
+GenieSim-Instruction focuses on language-conditioned manipulation, probing how well policies can follow diverse and compositional natural language instructions.
+GenieSim-Robust systematically evaluates policy robustness by introducing controlled perturbations across robot configuration, camera parameters, control latency, and environmental conditions.
+GenieSim-Manipulation covers a broad set of challenging manipulation tasks that require precise motor skills, long-horizon planning, and dexterous object interactions.
+Together, these four suites provide complementary perspectives on policy capabilities and benchmark reliability.
+Figure 6: GenieSim-Sim2Real Task Setups.
+Comparison of initial task configurations between real and simulated testing environments across all evaluated tasks.
+Rows 1 and 3 show real-world environments, while rows 2 and 4 show their corresponding simulated counterparts.
+V-A Experimental Setup
+GenieSim-Sim2Real.
+All experiments within this suite use the $\pi_{0.5}$ model [undefau] as the base policy, post-trained on varying amounts of real-world or synthetic data.
+The Agibot G1 robot is deployed for physical task execution.
+Each training configuration is evaluated over 50 trials in real-world environments and in simulation, ensuring direct comparability between the two evaluation modalities.
+GenieSim-Instruction, GenieSim-Robust, and GenieSim-Manipulation.
+For these three benchmark suites, three state-of-the-art vision-language-action models are compared: $\pi_{0.5}$ [undefau], GR00T-N1.6 [undefay], and $\pi_{0}$ [undefaz].
+Each model is fine-tuned with task-specific synthetic data generated by GenieSim and evaluated on the Agibot G2 robot using GenieSim’s automated closed-loop evaluation pipeline.
+V-B GenieSim-Sim2Real
+V-B1 Synthetic Data Scaling
+TABLE I: Synthetic Data Scaling Results. Task success rates of $\pi_{0.5}$ fine-tuned with varying amounts of real and synthetic data, evaluated in both simulated and real environments.
 Select Color Recognize Size Grasp Targets Organize Objects
@@
 500 eps sim 0.53 0.60 0.50 0.63 0.29 0.33 0.39 0.35
-1500 eps sim 0.86 0.85 0.93 0.94 0.80 0.71 0.52 0.60
-The validity and reliability of simulation benchmarks and datasets constitute a foundational consideration in embodied intelligence research.
-Figure 6: Sim to Real Experiments. Comparison of initial task setup between real and sim testing environments
-Two central questions are examined: first, whether a simulation benchmark can evaluate model performance as reliably as physical environment; and second, to what extent data synthetic data can effectively substitute for physically-collected data in model training.
-To address these questions systematically, a series of experiments are designed to assess the validity and efficacy of Genie Sim 3.0 benchmark and dataset.
-V-A Experimental Setup
-In this work, we conducted 32 groups of experiments to compare the task success rates of models fine-tuned with varying amounts of real-world and synthetic data, tested in both physics and simulation environments.
-The $\pi_{0.5}$ model [undefau]— a widely recognized baseline with demonstrated performance in related research— is selected as the base model and Agibot G1 robot is deployed for task execution.
-Since the primary objective of these experiments is to evaluate the validity and reliability of the simulation benchmark and dataset, tasks with either excessively high or low success rates are avoided to better reveal performance differences across dataset and environments.
-In accordance with this principle and to ensure assessment covering both atomic skills and cognitive capabilities, four representative tasks are selected for evaluation:
-  1. 1.
-    Select Color,
-  2. 2.
-    Recognize Size,
-  3. 3.
-    Grasp Targets,
-  4. 4.
-    Organize Items.
-For each task, models are post-trained with four different setups:
-  1. 1.
-    200 episodes of real-env data,
-  2. 2.
-    500 episodes of real-env data,
-  3. 3.
-    500 episodes of sim-env data,
-  4. 4.
-    1500 episodes of sim-env data.
-Each experiment is evaluated over 50 trials in real world environment and over 250 trails in simulation.
-Figure 7: Model Performance on Real-World Environment. The success rates of the 4 tasks, each with four training setups are shown in the figure with different colors.
-V-B Results and Analysis
-Figure 8: Performance Comparison between Real and Sim Environments. The performance in simulation share the same trend with it in real-world environment.
-Effectiveness of the Synthetic Data.
-In general, the success rates of the tasks significantly improve with an increasing volume of training data—for both real and synthetic data—which aligns well with the scaling laws in learning performance, see Fig.
-7.
-Increased task complexity leads to reduced success rates for models trained on datasets of identical size.
-Most importantly, the model trained exclusively on 1500 eps of synthetic data attains the highest zero‑shot success rates across all four tasks in the real‑world environment, as shown in Tab.
+1500 eps sim 0.86 0.85 0.93 0.94 0.72 0.71 0.52 0.60
+We conduct 32 groups of experiments across four representative tasks: Select Color, Recognize Size, Grasp Targets, and Organize Items.
+These tasks span a range of complexities—from atomic manipulation skills to language-conditioned cognitive reasoning—while avoiding floor or ceiling effects that would obscure performance differences.
+For each task, $\pi_{0.5}$ is fine-tuned under four data configurations: 200 and 500 episodes of real-world data, and 500 and 1500 episodes of synthetic data.
+Results are reported in Tab.
 I.
-When comparing data of the same quantity (500 eps), models trained on real‑robot data consistently outperform those trained on synthetic data.
-This advantage can be attributed to the higher physical fidelity of real data, which inherently captures complex real‑world phenomena such as friction, contact dynamics, and collisions—aspects that are often imperfectly modeled and lead to sim‑to‑real gaps in synthetic datasets.
-Figure 9: Correlation Analysis of Model Performance on Sim and Real Environments.
-All 16 models are evaluated both in sim and real environments, which are tagged on the right side of the figure, each with a different color and marker.
-Figure 10: Success Rate on both Sim and Real Env in Heatmap. The comparison of the left and right charts intuitively reflects the consistency between the real and the sim environment.
-Nevertheless, real‑world robotics data collection incurs substantial costs and time constraints, limiting the scale and diversity of available datasets.
-In contrast, synthetic data can be generated efficiently at scale, allowing for extensive randomization in parameters such as object textures, lighting conditions, physics properties, and task variations.
-By scaling the synthetic dataset to 1500 episodes and systematically varying simulation conditions, the resulting training distribution effectively mitigates the domain gap, thereby enhancing the model’s generalization capabilities to real‑world settings.
-Thus, the large‑scale, systematically randomized synthetic data pipeline implemented in GenieSim 3.0 offers a pragmatic and scalable alternative to real‑robot data collection, enabling the training of robust models capable of high zero‑shot performance in physical environments.
+Effectiveness of Synthetic Data.
+Success rates improve monotonically with increasing data volume for both real and synthetic sources, consistent with scaling laws in robot learning.
+At equivalent scale (500 eps), models trained on real-world data outperform those trained on synthetic data, reflecting the higher physical fidelity of real demonstrations.
+However, scaling synthetic data to 1500 episodes enables the model to surpass all real-data baselines across all four tasks in zero-shot real-world evaluation.
+This demonstrates that Genie Sim’s systematic domain randomization—spanning object textures, lighting conditions, physics parameters, and task variations—effectively bridges the domain gap at scale, offering a practical and scalable alternative to real-world data collection.
 Consistency between Sim and Real Environments.
-Fig.
-8 provides a comparative analysis of model performance in simulated versus physical environments.
-While minor discrepancies are observed across tasks—primarily attributable to systematic differences from approximations of contact dynamics and friction in simulation, as well as stochastic factors inherent to real-world deployment such as unstructured environmental perturbations—the overall performance trends remain consistent.
-This close correspondence is further illustrated in the success-rate heatmaps presented in the Fig.
-10.
-A quantitative correlation analysis, summarized in the scatter plot of the Fig.
-9, reveals a statistically significant strong positive relationship between simulated and real-world performance ($R^{2}=0.924$), with a best-fit slope $\approx 1.045$.
-These results indicate that evaluations conducted in Genie Sim 3.0 yield reliable conclusions that are consistent with trends observed in real-world testing.
-Given this high level of alignment, alongside the advantages of simulation—including fully automated evaluation pipelines, high reproducibility, scalable scenario generation, support for concurrent batch executions, and significantly lower operational costs—Genie Sim 3.0 serves as a robust and efficient platform for model assessment.
-Its use can substantially accelerate research iteration and facilitate experimental scaling beyond the practical constraints of physical testing.
+Model performance across simulated and physical environments is compared.
+Despite minor discrepancies attributable to imperfect simulation of contact dynamics and stochastic real-world perturbations, overall performance trends remain consistent across environments.
+A quantitative correlation analysis (Fig.
+7) yields $R^{2}=0.94$ with a slope $\approx 1.025$, indicating that evaluations of Genie Sim reliably predict real-world performance trends and can serve as a cost-efficient substitute for physical benchmarking.
+Figure 7: Correlation Analysis of Sim and Real Performance. All 16 model configurations show a strong linear relationship between simulated and real-world success rates ($R^{2}=0.94$).
+V-B2 Sim-to-Real Transfer
+[…]
+Tasks Sim Env Real Env
+Sim Data Real Data Sim Data Real Data
+(sim-to-sim) (real-to-sim) (sim-to-real) (real-to-real)
+Select Color 0.86 0.75 0.85 0.73
+Recognize Size 0.93 0.75 0.94 0.75
+Grasp Targets 0.72 0.54 0.71 0.58
+Organize Items 0.52 0.45 0.60 0.40
+Pack in Supermarket 0.94 1.00 0.95 0.95
+Sort Fruit 0.90 0.90 1.00 1.00
+Place Block into Drawer 0.80 0.90 0.85 0.90
+Bimanual Chip Handover 0.80 0.70 0.73 0.71
+Avg. 0.81 0.75 0.83 0.75
+$\dagger$ Sim Data: $500\sim 1500$ episodes of simulation data. Real Data: $500$ episodes of real-world data.
+To further validate the cross-domain transfer capability of Genie Sim, we extend the evaluation to 8 tasks in Tab.
+II, adding four newly introduced tasks—Pack in Supermarket, Sort Fruit, Place Block into Drawer, and Bimanual Chip Handover—to the original four.
+For each task, models trained on synthetic data (50050015001500 episodes) are compared against models trained on 500 episodes of real-world data, with both evaluated in simulated and physical environments.
+Sim-to-Real.
+Across all 8 tasks, models trained solely on synthetic data achieve competitive or superior real-world performance relative to models trained on real data, with an average sim-to-real success rate of 0.83 versus 0.75 for real-to-real.
+This confirms that synthetic data generated by Genie Sim is of sufficient fidelity for direct real-world deployment.
+Real-to-Sim.
+Conversely, models trained on real-world data also transfer well to simulated evaluation, achieving an average of 0.75.
+This bidirectional transferability demonstrates that the Genie Sim environment faithfully captures the physical structures and task dynamics of real-world settings, establishing it as a reliable evaluation platform.
+Taken together, the 32-group scaling experiments and the 8-task transfer study demonstrate that Genie Sim provides both high-quality synthetic data and a reliable simulation benchmark—the two foundational properties required for the credible cross-model evaluations conducted in the following suites.
+V-C Genie Sim Benchmark Evaluation
+Grounded in the validated reliability of Genie Sim, we present a systematic comparison of $\pi_{0.5}$, GR00T-N1.6, and $\pi_{0}$ across three complementary benchmark suites.
+Each suite targets a distinct capability dimension, and the high sim-to-real consistency established in Sec.
+V-B ensures that simulation-based rankings are predictive of real-world policy performance.
+GenieSim-Instruction.
+This suite comprises 10 tasks that require policies to interpret and execute diverse natural language instructions, including directives based on object number, shape, type, color, size, and commonsense or logical reasoning.
+The tasks are designed to stress-test the language grounding capabilities of VLA models beyond simple keyword matching.
+As shown in Tab.
+III, $\pi_{0.5}$ achieves an average success rate of 0.67, substantially outperforming GR00T-N1.6 (0.40) and $\pi_{0}$ (0.28).
+The performance gap widens on semantically complex tasks such as pick_common_sense (0.35 vs.
+0.12 vs.
+0.05), indicating that $\pi_{0.5}$’s stronger language model backbone provides a meaningful advantage in instruction-following scenarios.
+TABLE III: Experimental results on GenieSim-Instruction. Success rates across 10 language-conditioned manipulation tasks.
+Tasks $\boldsymbol{\pi_{0.5}}$ GR00T-N1.6 $\boldsymbol{\pi_{0}}$
+pick_block_number 0.73 0.28 0.17
+pick_block_shape 0.41 0.15 0.17
+pick_common_sense 0.35 0.12 0.05
+pick_follow_logic_or 0.58 0.56 0.26
+pick_object_type 0.81 0.56 0.27
+pick_specific_object 0.58 0.35 0.16
+straighten_object 0.66 0.33 0.46
+pick_billiards_color 0.81 0.37 0.47
+pick_block_color 0.88 0.71 0.40
+pick_block_size 0.89 0.52 0.36
+Avg. 0.67 0.40 0.28
+TABLE IV: Experimental results on GenieSim-Robust.
+Success rates under 12 categories of environmental and control perturbations, with $\Delta$ denoting the change relative to the Reference condition.
+The significant drops per model are shown in bold.
+Generalization $\boldsymbol{\pi_{0.5}}$ GR00T-N1.6 $\boldsymbol{\pi_{0}}$
+Score $\Delta$ Score $\Delta$ Score $\Delta$
+Reference 0.92 0.58 0.46
+Instruction 0.89 -0.03 0.47 -0.11 0.32 -0.14
+Robot Init Base 0.83 -0.09 0.59 +0.01 0.32 -0.14
+Robot Init Joint 0.70 -0.22 0.39 -0.19 0.34 -0.12
+Robot End Effector 0.42 -0.50 0.30 -0.28 0.26 -0.20
+Control Delay 0.76 -0.16 0.57 -0.01 0.40 -0.06
+Camera Frame Drop 0.83 -0.09 0.28 -0.30 0.19 -0.27
+Camera Noise 0.89 -0.03 0.59 -0.01 0.34 -0.12
+Camera Occlusion 0.93 +0.01 0.59 +0.01 0.41 -0.05
+Camera Extrinsic 0.39 -0.53 0.27 -0.31 0.22 -0.24
+Ambient Lighting 0.85 -0.07 0.54 -0.04 0.44 -0.02
+Background 0.90 -0.02 0.57 -0.01 0.40 -0.06
+Avg. 0.77 0.48 0.34
+TABLE V: Experimental results on GenieSim-Manipulation. Success rates across 10 complex manipulation tasks.
+Tasks $\pi_{0.5}$ GR00T-N1.6 $\pi_{0}$
+Open Door 0.60 0.35 0.46
+Hold Pot 0.35 0.00 0.14
+Pour Workpiece 0.95 0.95 0.72
+Stock and Straighten Shelf 0.37 0.15 0.21
+Take Wrong Item Shelf 0.95 0.65 0.80
+Scoop Popcorn 0.78 0.80 0.68
+Clean the Desktop 0.16 0.01 0.08
+Place Block into Box 0.50 0.30 0.38
+Sorting Packages 0.45 0.14 0.13
+Sorting Packages Continuous 0.16 0.03 0.00
+Avg. 0.53 0.34 0.36
+GenieSim-Robust.
+This suite systematically evaluates policy robustness by introducing 12 categories of controlled perturbations to a reference pick-and-place task, spanning robot initial pose, end-effector configuration, control delay, camera frame drop, image noise, occlusion, extrinsic calibration error, ambient lighting, and scene background.
+These conditions model realistic deployment variabilities that are difficult to enumerate exhaustively in physical experiments but are readily programmable in simulation.
+As reported in Tab.
+IV, $\pi_{0.5}$ achieves the highest average success rate of 0.77 across all perturbation categories, compared to 0.48 for GR00T-N1.6 and 0.34 for $\pi_{0}$.
+The relative degradation from the Reference condition is smallest for $\pi_{0.5}$, indicating that it not only achieves higher absolute performance but also degrades more gracefully under distribution shift.
+Looking at specific perturbation categories, $\pi_{0.5}$ exhibits the largest drops under Camera Extrinsic ($\Delta=-0.53$) and Robot End Effector ($\Delta=-0.50$) conditions, suggesting sensitivity to physical calibration errors and changes in end-effector configuration.
+Robot Init Joint also causes a notable decline ($\Delta=-0.22$).
+In contrast, $\pi_{0.5}$ shows strong robustness to perceptual perturbations such as Camera Occlusion ($\Delta=+0.01$), Camera Noise ($\Delta=-0.03$), and scene Background changes ($\Delta=-0.02$), indicating effective visual invariance to appearance-level disturbances.
+For GR00T-N1.6, the most significant performance drops occur under Camera Extrinsic ($\Delta=-0.31$), Camera Frame Drop ($\Delta=-0.30$), and Robot Init Joint ($\Delta=-0.19$), revealing particular vulnerability to temporal input degradation and geometric miscalibration.
+Meanwhile, GR00T-N1.6 maintains relatively stable performance under Control Delay ($\Delta=-0.01$), Camera Noise ($\Delta=-0.01$), and Background ($\Delta=-0.01$), demonstrating reasonable tolerance to low-level signal perturbations.
+$\pi_{0}$ suffers the greatest degradation from Camera Frame Drop ($\Delta=-0.27$) and Camera Extrinsic ($\Delta=-0.24$), followed by Robot End Effector ($\Delta=-0.20$), indicating that its visual perception and spatial reasoning are most susceptible to input incompleteness and calibration shifts.
+On the other hand, $\pi_{0}$ shows comparatively better resilience to Ambient Lighting ($\Delta=-0.02$) and Camera Occlusion ($\Delta=-0.05$), suggesting some degree of robustness to illumination and partial visual obstruction.
+Across all three models, Camera Extrinsic and Robot End Effector perturbations consistently induce the largest performance drops, pointing to these as the most challenging axes of distributional shift for current visuomotor policies.
+GenieSim-Manipulation.
+This suite consists of 10 tasks that target diverse and challenging manipulation skills, including articulated object interaction (Open Door), bimanual collaboration (Hold Pot), granular material handling (Scoop Popcorn), precise placement and sorting (Place Block into Box, Sorting Packages), desktop cleaning, and multi-step shelf organization.
+These tasks collectively probe the breadth and depth of a policy’s physical manipulation capabilities in realistic scenarios.
+As shown in Tab.
+V, $\pi_{0.5}$ achieves the highest average success rate of 0.53, ahead of $\pi_{0}$ (0.36) and GR00T-N1.6 (0.34).
+High-performance tasks such as Pour Workpiece and Take Wrong Item Shelf are handled competitively by all three models, reflecting the robustness of large pretrained VLA backbones on simpler manipulation primitives.
+In contrast, long-horizon sequential tasks—Sorting Packages Continuous (0.16 / 0.03 / 0.00) and Clean the Desktop (0.16 / 0.01 / 0.08)—remain challenging for all models, highlighting the frontier of current policy capabilities and the discriminative value of this suite.
+Across all three task suites, $\pi_{0.5}$ consistently achieves the highest average performance, demonstrating advantages in instruction following, perturbation robustness, and manipulation dexterity.
+The credibility of these cross-model comparisons rests on the reliability of Genie Sim as a benchmark platform, which has been rigorously validated through the sim-to-real experiments in Sec.
+V-B.
 VI CONCLUSIONS
@@
 ACKNOWLEDGMENT
-We would like to express our sincere gratitude to Sukai Wang and Xiaowei Cai for their assistance with model inference and training; to Haowen Yin, Yanping Zhou, Zengcheng Zhou, and Cheng Ruan for their support in experimental testing; to Chengyue Zhao, Xinke Yu, and Jialu Li for their work on publicity and design; and to Cheng Jing, Haoyu Cao, Chi Zhang, Guangte Xiang, Jun Dai for their help with data processing.
+We would like to express our sincere gratitude to Linqing Zhong, Sukai Wang and Xiaowei Cai for their assistance with model inference and training; to Haowen Yin, Yanping Zhou, Jiahao Yu, Zengcheng Zhou, and Cheng Ruan for their support with experiment testing; to Chengyue Zhao, Xinke Yu, and Jialu Li for their work on publicity and design; and to Cheng Jing, Haoyu Cao, Chi Zhang, Guangte Xiang, Jun Dai, Jia Zeng for their help with data processing.